OpenStudy (anonymous):

PLEASE HELP!!! 3sqrt45xcubed+2sqrt12+sqrt27-3sqrt20xcubed

OpenStudy (anonymous):

Can you mke this more clear using the equation button?

OpenStudy (anonymous):

\[5^{3} \] see it's down by draw and attachfile. You can make it more clear. like above.

OpenStudy (anonymous):

\[3\sqrt{45x ^{3}}+2\sqrt{12}+\sqrt{27}-3\sqrt{20x ^{3}}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

3sqrt(45)sqrt(x^3)+2sqrt(3)sqrt(4)+sqrt(9)*sqrt(3)-3sqrt(4)sqrt(5)sqrt(x^3) = 3sqrt(9)*sqrt(5)*sqrt(x^3)+4sqrt(3)+3sqrt(3)-6sqrt(5)sqrt(x^3)

OpenStudy (zzr0ck3r):

with me so far?

OpenStudy (zzr0ck3r):

9sqrt(5x^3)-6sqrt(5x^3)+7sqrt(3) = 3sqrt(5x^3)+7sqrt(3)

OpenStudy (radar):